# Abundance of cosmic voids in EFT of dark energy

### v1

@@ I Introduction
Scalar-tensor theories constitute not only the simplest but also the most extensively studied class of modified gravity models, and can effectively describe a wide variety of modified gravity scenarios. In scalar-tensor theories, general relativity is extended by introducing an additional scalar degree of freedom on top of the two tensor degrees of freedom corresponding to gravitational waves. The Horndeski theory is the most general scalar-tensor theory with second-order field equations [Horndeski:1974wa, Deffayet:2011gz, Kobayashi:2011nu], and can describe each specific model of modified gravity when the functions in the action are specified. See Ref. [Kobayashi:2019hrl] for a review. It has been applied extensively to the study of cosmology in modified gravity (see, e.g., Ref. [Arai:2022zzz]).
@@
In the Horndeski family of theories, the characteristic screening mechanism, known as the Vainshtein mechanism, can operate due to nonlinear derivative interactions of the scalar degree of freedom, restoring the gravitational field predicted in general relativity in the vicinity of a source [Kimura:2011dc, Narikawa:2013pjr, Koyama:2013paa]. In the context of cosmology, the operation of this mechanism relies on the nonlinearity of matter density perturbations. In this paper, we focus on the large-scale structure (LSS) of the Universe, where the nonlinearity of matter density perturbations must be taken care of appropriately.
@@
The LSS of the Universe provides one of the most powerful observational probes for testing gravity on cosmological scales. It is composed of several elements, including halos, filaments, walls, and voids. Among these, we focus on cosmic voids, vast underdense regions that occupy most of the volume of the Universe. In contrast to halos, the density contrast $\delta$ in voids has the limit $\delta\geq-1$ by definition, and hence $|\delta|$ cannot be much larger than one. As a consequence, the Vainshtein screening mechanism is expected to work only partially in void environments, potentially leaving some signatures of modified gravity. Indeed, Ref. [Falck:2014jwa] demonstrated that particles inside voids feel unscreened scalar-mediated forces in the nDGP model, a modified gravity model that can be described effectively by a subset of the Horndeski family. For this reason, it is possible that cosmic voids serve as excellent laboratories for testing gravity on cosmological scales.
@@
Several studies have investigated the properties of cosmic voids in modified gravity. The authors of Ref. [2013MNRAS.431..749C] examined the evolution of voids in chameleon theories using an extended spherical collapse model based on the dynamics of spherical shells. In Ref. [Voivodic:2016kog], the abundance of cosmic voids was computed and compared with results from $N$-body simulations in $f(R)$ gravity and symmetron models. Building on simulation-based approaches, Ref. [2012MNRAS.421.3481L] analyzed the void population in $f(R)$ gravity with $N$-body simulations and computed the void number density as a function of volume. In the context of the nDGP model, Ref. [Falck:2017rvl] explored the impact of modified gravity on the density, velocity, and screening profiles of voids. A similar analysis was performed in Ref. [Perico:2019obq] for $f(R)$ gravity and symmetron models (see also Refs. [Wilson:2022ets, Winther:2015pta, Zivick:2014uva, Contarini:2020fdu, Cai:2014fma, Mauland:2023eax]). Weak-lensing signatures of voids were investigated in the nDGP model [Paillas:2018wxs, Davies:2019yif] and in the Galileon theories [Su:2022yoj, Baker:2018mnu, Barreira:2015vra].
@@
The studies mentioned above rely on specific models of modified gravity, each of which constitutes a subset of the Horndeski family. To investigate gravity theories in a more model-independent manner, in this work we employ the effective field theory (EFT) of dark energy, a general and unified framework for studying cosmology in scalar-tensor theories [Gubitosi:2012hu, Gleyzes:2013ooa, Bloomfield:2012ff, Bellini:2014fua, Cusin:2017mzw]. This approach enables us to analyze modifications of gravity without specifying the explicit functional forms of the free functions in the Horndeski action, while assuming the background evolution and the time dependence of the coefficient of each term in the effective action.
@@
In this paper, we model the void formation as the evolution of spherical mass shells. In this simplified modeling, the criterion for the formation of a void is given by the intersection of two neighboring shells at the outer boundary of the underdense region. One can then assess how the EFT of dark energy parameters change the linearly extrapolated critical density contrast for the void formation. Combining this with several other quantities affected by gravity modification, we calculate the void size function (VSF) [Sheth:2003py], i.e., the number density of voids as a function of their volume, and discuss the impacts of the EFT of dark energy parameters on it.
@@ II.2 Gravitational field equations in EFT of dark energy
The effective action for the metric in the Newtonian gauge Eq.(1), and for the scalar field, $\phi=t+\pi(t,\vec{x})$, is given by [Dima:2017pwp] (see also [Crisostomi:2017lbg, Langlois:2017dyl])
@@
If one starts from DHOST theories beyond Horndeski [Langlois:2015cwa], more terms appear in the effective action, accompanied by additional $\alpha$ functions. However, in this paper, we restrict ourselves to manifestly second-order scalar-tensor theories and use the EFT action (13) derived from the most general second-order scalar-tensor theories. We still have the four independent functions of time, $\alpha_{T}$, $\alpha_{B}$, $\alpha_{V}$, and $\alpha_{M}$. By requiring that the speed of gravitational waves is equal to that of light, we set
@@
Thus, only a single function $\alpha_{B}$ is left, which characterizes the kinetic gravity braiding effect that is the characteristic feature of the cubic galileon model [Nicolis:2008in, Deffayet:2009wt, Deffayet:2010qz, Kobayashi:2010cm]. In this setup, the coefficient of the nonlinear derivative interaction is characterized by $\alpha_{B}$.
@@ II.2.1 Background model
where $p$ is some constant and $\alpha_{i0}$ is the present value of $\alpha_{i}$. The second equality is valid from the matter-dominated era to the present time. The model with $p=1$ is most frequently used in the literature. The model with general $p$ was considered in Ref. [Traykova:2021hbr], in which it was reported that this parametrization fits more than $98\%$ of randomly generated shift-symmetric models with constant $G_{4}$ in the Horndeski action.
@@ II.2.2 Gravitational-field equations and equation of motion for the shells
For given $H(t)$, $\Omega_{\mathrm{m}}(t)$, and $\alpha_{B}(t)$, this equation can be used to determine the trajectory of a particle $R(t)$ in the EFT of dark energy. It is worth noting that the expression in the square root, $\alpha_{*}^{2}+2\alpha_{B}^{2}\delta m$, can be negative in an underdense region, $\delta m<0$, depending on the assumed cosmological background model. This would result in imaginary $\pi$, and hence we exclude such cosmological background models. The same problem has been pointed out already in Refs. [Winther:2015pta] in the context of cubic Galileon models.
@@ III.1 Initial profile of the underdense region and shell-crossing
To determine the moment of void formation, we first need to specify the condition that signals the formation of a void. A common criterion is that a void forms when shell-crossing occurs, i.e., two neighboring shells intersect. The moment of shell-crossing is determined from the condition [Moretti:2025gbp]
@@
where $R_{i}$ is the initial physical radius of the inner one of the two shells and $R(t,R_{i})$ is the physical radius of that shell at the time $t$. In general, the time $t=t_{*}$ satisfying Eq. (44) depends on $R_{i}$: $t_{*}(R_{i})$. Following Ref. [Moretti:2025gbp], we adopt the definition that the void forms at the moment $t=t_{\mathrm{sc}}$ when Eq. (44) is satisfied for the outermost shells (i.e., $t_{\mathrm{sc}}:=t_{*}(R_{0}+\Delta R)$). As shown in Fig. 1, for $s\gg 1$, the two outermost shells intersect first for our choice of the initial conditions specified below. With this definition of $t_{\mathrm{sc}}$, the known analytical result in the Einstein-de Sitter universe, $\delta_{\mathrm{sc}}\simeq-2.718$, is reproduced when $s\gg 1$ is taken, which supports adopting the above definition in a more general setup [Moretti:2025gbp].
@@ III.2.1 Underdensity
Solving the evolution equation (33) for $R$ with the initial conditions (45), we obtain the shell-crossing time $t_{\mathrm{sc}}$ for different assumptions on the time dependence of $\alpha_{B}$ (i.e., the models with $p=1$ and $p=0.85$).
@@
where $A_{Li}=-\delta_{0}/a(t_{i})$. The resultant $\delta_{\mathrm{sc}}$ is shown in Fig. 2 as a function of the redshift $z=a^{-1}(t_{\mathrm{sc}})-1$ at shell-crossing. In both $p=1$ and $p=0.85$ models, it can be seen that the critical density contrast increases as $\alpha_{B0}$ increases when compared at the same redshift. This can be understood as follows: as $\alpha_{B0}$ increases, gravity becomes stronger and the shell-crossing occurs earlier, which tends to make $|\delta_{\mathrm{sc}}|$ smaller. In contrast, the linear growth of $\delta_{L}$ is enhanced with increasing $\alpha_{B0}$, which acts to make $|\delta_{\mathrm{sc}}|$ larger. The final value of $\delta_{\mathrm{sc}}$ is determined by the competition between these two effects. It is found that the two effects are comparable and nearly cancel each other out. As a result, the modification of $\delta_{\mathrm{sc}}$ is one order of magnitude smaller than the order of $\alpha_{B0}$. Since we have $\Omega_{\mathrm{m}}\to 1$ and $\alpha_{B}\to 0$ in the deep matter-dominated era, $\delta_{\mathrm{sc}}$ approaches the Einstein-de Sitter value at early times.
@@
which quantifies how much the comoving radius of a shell expands by the shell-crossing time. Figure 3 shows the numerical results for the expansion factor $\mathcal{R}$ as a function of redshift $z$, for the models with $p=1$ and $p=0.85$. We find that the expansion factor decreases as $\alpha_{B0}$ increases. This is because a larger value of $\alpha_{B0}$ strengthens gravity, thereby slowing down the expansion of each shell. Consequently, the comoving radius of the shell at the shell-crossing time becomes smaller.
One might be interested in how $\delta_{\mathrm{sc}}$ is modified in models with smaller values of $p$. However, we find that it is not possible to compute the shell evolution for $p\lesssim 0.85$ with the fiducial value $\alpha_{B0}=0.1$, because the term inside the square root in Eq. (33) becomes negative at some moment, causing the physical radius of the shell to become imaginary before the shell-crossing occurs for any initial amplitude $A_{Li}$ of $\delta$. Note that this occurs because we are considering an underdense region ($\delta m<0$).
@@ III.2.2 Overdensity
For the nonlinear evolution of a top-hat overdensity, we solve Eq. (86) with the initial conditions $\delta(t_{i})=A_{i}a(t_{i})$ and $\dot{\delta}(t_{i})=H(t_{i})\delta(t_{i})$, where $A_{i}$ is a constant. These initial conditions are justified because the linear evolution of $\delta$ is governed by the standard equation in the deep matter-dominated era. We suppose that the overdense region collapses when $\delta$ reaches $10^{7}$, and evaluate the collapse time $t_{\mathrm{col}}$ using the condition $\delta(t_{\mathrm{col}})=10^{7}$. Then, we calculate the critical density contrast $\delta_{\mathrm{c}}$ from the linear evolution, $\delta_{\mathrm{c}}=\delta_{L}(t_{\mathrm{col}})$. The resultant critical density contrast $\delta_{\mathrm{c}}$ is shown in Fig. 4 as a function of the redshift $z=a^{-1}(t_{\mathrm{col}})-1$ for the background model with $p=1$ and $p=0.85$.
@@ IV Implications for the void size function
To compare theoretical predictions with observational data of cosmic voids, it is crucial to study the VSF. To this end, we adopt the VSF proposed in Ref. [Sheth:2003py],
@@
and the linear matter power spectrum at redshift $z$, $P(k,z)$. The explicit form of the function $f(\sigma,\delta_{\mathrm{c}},\delta_{\mathrm{sc}})$ will be given below. The quantities $\delta_{\mathrm{sc}}$, $\mathcal{R}$, and $\delta_{\mathrm{c}}$ are computed in the way described in Sec. III.2. The matter power spectrum is computed using a modified version of the publicly available Einstein–Boltzmann solver hi_class [Zumalacarregui:2016pph]. When computing the power spectrum, we adopt the fiducial value $\alpha_{K0}=1$, unless stated otherwise. We are thus ready to calculate the VSF.
@@
The explicit form of function $f(\sigma,\delta_{\mathrm{c}},\delta_{\mathrm{sc}})$ is given by [Sheth:2003py, Jennings:2013nsa]
@@
It is not easy to compute the above infinite sum. However, in Ref. [Jennings:2013nsa], the authors showed that if $\mathcal{D}$ is smaller than $3/4$, this function can be approximated as
@@
Next, to disentangle the individual effects of the modification of gravity and to identify which component is responsible for the changes in the VSF, we compute the VSF by turning on each of the components $(\sigma,\delta_{\mathrm{c}},\delta_{\mathrm{sc}},\mathcal{R})$ one at a time. The results are presented in Fig. 6. From this figure, two features can be noted. On small scales, the change in the VSF comes mainly from the modification in the linear matter power spectrum. In contrast, on large scales, the modifications in $\sigma$ and $\delta_{\mathrm{sc}}$ together account for the change in the VSF, though the modified linear matter power spectrum still has the largest impact. On all scales, the modifications in $\mathcal{R}$ and $\delta_{c}$ have negligible impacts on the VSF.

### v2

@@ I Introduction
Scalar-tensor theories constitute not only the simplest but also the most extensively studied class of modified gravity models, and can effectively describe a wide variety of modified gravity scenarios. In scalar-tensor theories, general relativity is extended by introducing an additional scalar degree of freedom on top of the two tensor degrees of freedom corresponding to gravitational waves. The Horndeski theory is the most general scalar-tensor theory with second-order field equations [25, 17, 29], and can describe each specific model of modified gravity when the functions in the action are specified. See Ref. [30] for a review. It has been applied extensively to the study of cosmology in modified gravity (see, e.g., Ref. [2]).
@@
In the Horndeski family of theories, the characteristic screening mechanism, known as the Vainshtein mechanism, can operate due to nonlinear derivative interactions of the scalar degree of freedom, restoring the gravitational field predicted in general relativity in the vicinity of a source [27, 38, 31]. In the context of cosmology, the operation of this mechanism relies on the nonlinearity of matter density perturbations. In this paper, we focus on the large-scale structure (LSS) of the Universe, where the nonlinearity of matter density perturbations must be taken care of appropriately.
@@
The LSS of the Universe provides one of the most powerful observational probes for testing gravity on cosmological scales. It is composed of several elements, including halos, filaments, walls, and voids. Among these, we focus on cosmic voids, vast underdense regions that occupy most of the volume of the Universe. In contrast to halos, the density contrast $\delta$ in voids has the limit $\delta\geq-1$ by definition, and hence $|\delta|$ cannot be much larger than one. As a consequence, the Vainshtein screening mechanism is expected to work only partially in void environments, potentially leaving some signatures of modified gravity. Indeed, Ref. [21] demonstrated that particles inside voids feel unscreened scalar-mediated forces in the nDGP model, a modified gravity model that can be described effectively by a subset of the Horndeski family. For this reason, it is possible that cosmic voids serve as excellent laboratories for testing gravity on cosmological scales.
@@
Several studies have investigated the properties of cosmic voids in modified gravity. The authors of Ref. [11] examined the evolution of voids in chameleon theories using an extended spherical collapse model based on the dynamics of spherical shells. In Ref. [48], the abundance of cosmic voids was computed and compared with results from $N$-body simulations in $f(R)$ gravity and symmetron models. Building on simulation-based approaches, Ref. [35] analyzed the void population in $f(R)$ gravity with $N$-body simulations and computed the void number density as a function of volume. In the context of the nDGP model, Ref. [20] explored the impact of modified gravity on the density, velocity, and screening profiles of voids. A similar analysis was performed in Ref. [41] for $f(R)$ gravity and symmetron models (see also Refs. [49, 50, 52, 12, 10, 36]). Weak-lensing signatures of voids were investigated in the nDGP model [40, 15] and in the Galileon theories [45, 3, 4].
@@
The studies mentioned above rely on specific models of modified gravity, each of which constitutes a subset of the Horndeski family. To investigate gravity theories in a more model-independent manner, in this work we employ the effective field theory (EFT) of dark energy, a general and unified framework for studying cosmology in scalar-tensor theories [24, 22, 8, 7, 14]. This approach enables us to analyze modifications of gravity without specifying the explicit functional forms of the free functions in the Horndeski action, while assuming the background evolution and the time dependence of the coefficient of each term in the effective action.
@@
In this paper, we model the void formation as the evolution of spherical mass shells. In this simplified modeling, the criterion for the formation of a void is given by the intersection of two neighboring shells at the outer boundary of the underdense region. One can then assess how the EFT of dark energy parameters change the linearly extrapolated critical density contrast for the void formation. Combining this with several other quantities affected by gravity modification, we calculate the void size function (VSF) [44], i.e., the number density of voids as a function of their volume, and discuss the impacts of the EFT of dark energy parameters on it.
@@ II.2 Gravitational field equations in EFT of dark energy
The effective action for the metric in the Newtonian gauge Eq.(1), and for the scalar field, $\phi=t+\pi(t,\vec{x})$, is given by [19] (see also [13, 34])
@@
If one starts from DHOST theories beyond Horndeski [33], more terms appear in the effective action, accompanied by additional $\alpha$ functions. However, in this paper, we restrict ourselves to manifestly second-order scalar-tensor theories and use the EFT action (13) derived from the most general second-order scalar-tensor theories. We still have the four independent functions of time, $\alpha_{T}$, $\alpha_{B}$, $\alpha_{V}$, and $\alpha_{M}$. By requiring that the speed of gravitational waves is equal to that of light, we set
@@
Thus, only a single function $\alpha_{B}$ is left, which characterizes the kinetic gravity braiding effect that is the characteristic feature of the cubic galileon model [39, 16, 18, 28]. In this setup, the coefficient of the nonlinear derivative interaction is characterized by $\alpha_{B}$.
@@ II.2.1 Background model
where $p$ is some constant and $\alpha_{i0}$ is the present value of $\alpha_{i}$. The second equality is valid from the matter-dominated era to the present time. The model with $p=1$ is most frequently used in the literature. The model with general $p$ was considered in Ref. [47], in which it was reported that this parametrization fits more than $98\%$ of randomly generated shift-symmetric models with constant $G_{4}$ in the Horndeski action.
@@ II.2.2 Gravitational-field equations and equation of motion for the shells
For given $H(t)$, $\Omega_{\mathrm{m}}(t)$, and $\alpha_{B}(t)$, this equation can be used to determine the trajectory of a particle $R(t)$ in the EFT of dark energy. It is worth noting that the expression in the square root, $\alpha_{*}^{2}+2\alpha_{B}^{2}\delta m$, can be negative in an underdense region, $\delta m<0$, depending on the assumed cosmological background model. This would result in imaginary $\pi$, and hence we exclude such cosmological background models. The same problem has been pointed out already in Refs. [50] in the context of cubic Galileon models.
@@ III.1 Initial profile of the underdense region and shell-crossing
To determine the moment of void formation, we first need to specify the condition that signals the formation of a void. A common criterion is that a void forms when shell-crossing occurs, i.e., two neighboring shells intersect. The moment of shell-crossing is determined from the condition [37]
@@
where $R_{i}$ is the initial physical radius of the inner one of the two shells and $R(t,R_{i})$ is the physical radius of that shell at the time $t$. In general, the time $t=t_{*}$ satisfying Eq. (44) depends on $R_{i}$: $t_{*}(R_{i})$. Following Ref. [37], we adopt the definition that the void forms at the moment $t=t_{\mathrm{sc}}$ when Eq. (44) is satisfied for the outermost shells (i.e., $t_{\mathrm{sc}}:=t_{*}(R_{0}+\Delta R)$). As shown in Fig. 1, for $s\gg 1$, the two outermost shells intersect first for our choice of the initial conditions specified below. With this definition of $t_{\mathrm{sc}}$, the known analytical result in the Einstein-de Sitter universe, $\delta_{\mathrm{sc}}\simeq-2.718$, is reproduced when $s\gg 1$ is taken, which supports adopting the above definition in a more general setup [37].
@@ III.2.1 Underdensity
Solving the evolution equation (33) for $R$ with the initial conditions (45), we obtain the shell-crossing time $t_{\mathrm{sc}}$ for different assumptions on the time dependence of $\alpha_{B}$ (i.e., the models with $p=1$ and $p=0.87$).
@@
where $A_{Li}=-\delta_{0}/a(t_{i})$. The resultant $\delta_{\mathrm{sc}}$ is shown in Fig. 2 as a function of the redshift $z=a^{-1}(t_{\mathrm{sc}})-1$ at shell-crossing. In both $p=1$ and $p=0.87$ models, it can be seen that the critical density contrast increases as $\alpha_{B0}$ increases when compared at the same redshift. This can be understood as follows: as $\alpha_{B0}$ increases, gravity becomes stronger and the shell-crossing occurs earlier, which tends to make $|\delta_{\mathrm{sc}}|$ smaller. In contrast, the linear growth of $\delta_{L}$ is enhanced with increasing $\alpha_{B0}$, which acts to make $|\delta_{\mathrm{sc}}|$ larger. The final value of $\delta_{\mathrm{sc}}$ is determined by the competition between these two effects. It is found that the two effects are comparable and nearly cancel each other out. As a result, the modification of $\delta_{\mathrm{sc}}$ is one order of magnitude smaller than the order of $\alpha_{B0}$. Since we have $\Omega_{\mathrm{m}}\to 1$ and $\alpha_{B}\to 0$ in the deep matter-dominated era, $\delta_{\mathrm{sc}}$ approaches the Einstein-de Sitter value at early times.
@@
which quantifies how much the comoving radius of a shell expands by the shell-crossing time. Figure 3 shows the numerical results for the expansion factor $\mathcal{R}$ as a function of redshift $z$, for the models with $p=1$ and $p=0.87$. We find that the expansion factor decreases as $\alpha_{B0}$ increases. This is because a larger value of $\alpha_{B0}$ strengthens gravity, thereby slowing down the expansion of each shell. Consequently, the comoving radius of the shell at the shell-crossing time becomes smaller.
One might be interested in how $\delta_{\mathrm{sc}}$ is modified in models with smaller values of $p$. However, we find that it is not possible to compute the shell evolution for $p\lesssim 0.87$ with the fiducial value $\alpha_{B0}=0.1$, because the term inside the square root in Eq. (33) becomes negative at some moment, causing the physical radius of the shell to become imaginary before the shell-crossing occurs for any initial amplitude $A_{Li}$ of $\delta$. Note that this occurs because we are considering an underdense region ($\delta m<0$).
@@ III.2.2 Overdensity
For the nonlinear evolution of a top-hat overdensity, we solve Eq. (86) with the initial conditions $\delta(t_{i})=A_{i}a(t_{i})$ and $\dot{\delta}(t_{i})=H(t_{i})\delta(t_{i})$, where $A_{i}$ is a constant. These initial conditions are justified because the linear evolution of $\delta$ is governed by the standard equation in the deep matter-dominated era. We suppose that the overdense region collapses when $\delta$ reaches $10^{7}$, and evaluate the collapse time $t_{\mathrm{col}}$ using the condition $\delta(t_{\mathrm{col}})=10^{7}$. Then, we calculate the critical density contrast $\delta_{\mathrm{c}}$ from the linear evolution, $\delta_{\mathrm{c}}=\delta_{L}(t_{\mathrm{col}})$. The resultant critical density contrast $\delta_{\mathrm{c}}$ is shown in Fig. 4 as a function of the redshift $z=a^{-1}(t_{\mathrm{col}})-1$ for the background model with $p=1$ and $p=0.87$.
@@ IV Implications for the void size function
To compare theoretical predictions with observational data of cosmic voids, it is crucial to study the VSF. To this end, we adopt the VSF proposed in Ref. [44],
@@
and the linear matter power spectrum at redshift $z$, $P(k,z)$. The explicit form of the function $f(\sigma,\delta_{\mathrm{c}},\delta_{\mathrm{sc}})$ will be given below. The quantities $\delta_{\mathrm{sc}}$, $\mathcal{R}$, and $\delta_{\mathrm{c}}$ are computed in the way described in Sec. III.2. The matter power spectrum is computed using a modified version of the publicly available Einstein–Boltzmann solver hi_class [53]. When computing the power spectrum, we adopt the fiducial value $\alpha_{K0}=1$, unless stated otherwise. We are thus ready to calculate the VSF.
@@
The explicit form of function $f(\sigma,\delta_{\mathrm{c}},\delta_{\mathrm{sc}})$ is given by [44, 26]
@@
It is not easy to compute the above infinite sum. However, in Ref. [26], the authors showed that if $\mathcal{D}$ is smaller than $3/4$, this function can be approximated as
@@
Next, to disentangle the individual effects of the modification of gravity and to identify which component is responsible for the changes in the VSF, we compute the VSF by turning on each of the components $(\sigma,\delta_{\mathrm{c}},\delta_{\mathrm{sc}},\mathcal{R})$ one at a time. The results are presented in Fig. 6. From this figure, two features can be noted. On small scales, the change in the VSF comes mainly from the modification in the linear matter power spectrum, since the changes that come from $\delta_{\mathrm{c}}$ and $\delta_{\mathrm{sc}}$ compensate. In contrast, on large scales, the modifications in $\sigma$ and $\delta_{\mathrm{sc}}$ together account for the change in the VSF, though the modified linear matter power spectrum still has the largest impact. On all scales, the modifications in $\mathcal{R}$ and $\delta_{c}$ have negligible impacts on the VSF.